\begin{thebibliography}{1}

\bibitem{IndManWar}
Karl-Heinz Indlekofer, Eugenijus Manstavicius, and Richard Warlimont.
\newblock On a certain class of infinite products and arithmetical semigroups.
\newblock {\em Archiv der Mathematik}, 56:446--453, 1991.

\bibitem{IndWeh}
Karl-Heinz Indlekofer and Stefan Wehmeier.
\newblock Mean behaviour of multiplicative functions on additive arithmetical
  semigroups.
\newblock Submitted to Computers and Mathematics With Applications, 2004.

\bibitem{Kn1}
John Knopfmacher.
\newblock {\em Abstract analytic number theory}.
\newblock North Holland publishing company, 1975.

\bibitem{Kn2}
John Knopfmacher.
\newblock {\em Analytic arithmetic of algebraic function fields}.
\newblock Marcel Dekker, 1979.

\bibitem{Kn3}
John Knopfmacher and Wen-Bin Zhang.
\newblock {\em Number theory arising from finite fields. Analytic and
  probabilistic theory}, volume 241 of {\em Pure and applied mathematics}.
\newblock Marcel Dekker, New York, 2001.

\bibitem{LindqvistPeetre2}
Peter Lindqvist and Jaak Peetre.
\newblock On a number theoretic sum considered by {M}eissel --- a historical
  observation.
\newblock {\em Nieuw Arch. Wisk.}, 15(3):175--179, 1997.

\bibitem{Mertens}
Franz Mertens.
\newblock Ein {B}eitrag zur analytischen {Z}ahlentheorie.
\newblock {\em Journal f. reine und angewandte Mathematik}, 78:46--62, 1874.

\end{thebibliography}

\documentclass[12pt, a4paper]{article}
\usepackage{amsmath}
\usepackage{amsthm}
\usepackage{amsfonts}
\usepackage{amssymb}
\usepackage{latexsym}

\newcommand{\N}{\ensuremath{\mathbb N}}

\renewcommand{\P}{{\mathbb P}}

\newcommand{\A}{axiom ${\cal A}$ }

\renewcommand{\deg}{\partial}

\theoremstyle{definition}
\newtheorem{df}{Def.}[section]

\theoremstyle{plain}
\newtheorem{sa}[df]{Theorem}
\newtheorem{lem}[df]{Lemma}
\newtheorem{prop}[df]{Proposition}
\newtheorem{cor}[df]{Corollary}

\title{{\bf Meissel's theorem in additive arithmetical semigroups}}
\author{{\bf Stefan Wehmeier}\\ {\small Universit\"at Paderborn, Germany}}
\date{}

\begin{document}
\maketitle

\begin{abstract}
We show how to control the error term in Mertens' formula and related theorems in the context of additive arithmetical semigroups and carry over an old related result of Meissel. 

\smallskip 
2000 Mathematics Subject Classification: 11N80, 11A41, 11N05 

Keywords: arithmetical semigroups, Mertens' theorem
\end{abstract}

\section{Introduction}

Many algebraic and combinatorial structures may be viewed as multisets of indecomposable components. As a general framework, John Knopfmacher defined the notion of \emph{additive arithmetical semigroup} in his influential books \cite{Kn1} and \cite{Kn2}.

\smallskip

An additive arithmetical semigroup is a free commutative monoid $G$ with non-empty set of generators $P$, equipped with a non-negative integer-valued degree mapping $\deg$ such that
\[ \deg(ab) = \deg(a) + \deg(b) \] 
for all $a, b \in G$, and such that 
\[ G(n) := \# \{ a \in G; \deg(a) = n \} < \infty \] 
for all $n \in \N_0$. Equivalently, we may require that
\[ P(n) := \# \{ p \in P; \deg(p) = n \} \]
be finite for all $n$. In particular, $G$ is countable.

\smallskip

We denote the generating function by
\[ Z(y) := \sum_{n=0}^{\infty} G(n)y^n . \]
Its radius of convergence $\rho$ must satisfy $0 \leq \rho \leq 1$ since $\# G = \sum_n G(n)$ is infinite. Any sensible classification must treat the cases $\rho=0$ and $\rho=1$ separately; in this paper, we restrict our attention to the case $0 < \rho < 1$. Following the traditional notation by Knopfmacher, we set
\[ q:= 1/\rho . \]

\medskip

In the past decades, many authors have investigated the connection between $P(n)$ and $G(n)$. 
A usual hypothesis is that 
\begin{equation}
\label{AxiomA}
\frac{G(n)}{q^n} = A + r(n) 
\end{equation}
 with some $A > 0$ and
\[ r(n) \rightarrow 0 \] sufficiently fast; we term this condition \A (with error term $r(n)$). A desired result would be to show that 
\begin{equation}
\label{andef}
\lambda_n := \frac{n P(n)}{q^n} 
\end{equation}
converges to $1$, in analogy to the prime number theorem; however, this does not necessarily hold even if $r(n) \rightarrow 0$ exponentially fast, see \cite{IndManWar}.    
Wen-Bin Zhang has proved that if $\sum_{n=0}^{\infty} \sup_{k \geq n} |r(k)|$ converges, then $\lambda_n$ is bounded.
On the other hand, in a separate paper (\cite{IndWeh}), Indlekofer and I have shown that examples exist where $r_n = O(\ln(n)^2/n)$ and $\lambda_n$ is unbounded.

However, it is possible to show that the Cesaro limit of $\lambda_n$ exists; more precisely, the book by Zhang and Knopfmacher contains the following theorem (\cite{Kn3}, Section 3.3):
\begin{prop}
\label{TheoremErrorTerm}
Suppose that $\sum_{n=0}^{\infty} \sup_{k \geq n} |r(k)|$ converges.
Then, as $n \rightarrow \infty$,
\begin{itemize}
\item
\begin{equation}
\label{Pkkqkestimate}
 \sum_{k=1}^n \lambda_k = n + O(1). 
\end{equation}
\item 
\begin{equation}
\label{Pkqkestimate2}
 \sum_{k=1}^n \frac{P(k)}{q^k} = \ln(n) + C_1 + O(1/n)
\end{equation}
for some constant $C_1$.
\item
\begin{equation}
\label{ZhangMertens}
\prod_{k=1}^n \left(1 - 1/q^{k} \right)^{P(k)} = \frac{C_2}{n} +
O(1/n^2)
\end{equation}
for some constant $C_2$. 
\end{itemize} 
\end{prop}

In \cite{IndWeh}, Indlekofer and I have shown that Zhang's estimates hold under a much weaker condition, at the price of obtaining no error term:

\begin{prop}
\label{IndWehProp}
Suppose that
\begin{equation}
\label{Hycond}
Z(y) = \frac{H(y)}{1 - qy} 
\end{equation}
and $H(y) \rightarrow A>0$ as $y \rightarrow 1/q$.
Then, as $n \rightarrow \infty$,
\begin{itemize}
\item
\begin{equation}
\label{Lambdakqkestimate}
\sum_{\substack{p \in P, k \in \N \\ \deg(p^k) \leq n}} \frac{1}{k q^{k \deg(p)}} = \ln(n) + \gamma + \ln(A) + o(1)
\end{equation}
where $\gamma$ denotes Euler's constant.
\item
\begin{equation}
\label{Pkqkestimate}
\sum_{k=1}^n \frac{P(k)}{k} = \ln(n) + \gamma + \ln(A) - C_M + o(1) 
\end{equation}
where
\[ C_M = \sum_{k=1}^{\infty} P(k) \left(\frac{1}{q^k} - \ln\left(1 - \frac{1}{q^k}\right)\right) .\]
\item
\[ \sum_{k=1}^n \lambda_k = n + o(n)  . \]
\item
\begin{equation}
\label{Mertenseq}
 \prod_{k=1}^n \left(\frac{1}{1- 1/q^{k}}\right)^{P(k)} = A \exp(\gamma) n + o(n) . 
\end{equation}
\end{itemize}
\end{prop}
In particular, this gives us the values $C_1 = \gamma + \ln(A) - C_M$ and $C_2 = \frac{1}{A \exp(\gamma)}$ for the constants in (\ref{Pkqkestimate2}) and (\ref{ZhangMertens}). Since the proof is based on a Tauberian theorem, it is difficult to turn this into a theorem with explicit error terms.

\section{Mertens-type estimates error terms}

In this section, it will become apparent how the trade-off between hypotheses and error terms in Propositions \ref{TheoremErrorTerm} and \ref{IndWehProp} can be controlled.

\begin{lem}
\label{PkkGquotestimate}
Suppose \A. 
Then
\[ \sum_{k \in \N} k P(k) \frac{G(n-k)}{G(n)} = n - C_3 + o(1), \]
where
\begin{eqnarray*}
 C_3 & := &  \sum_{k=1}^{\infty} k P(k) \sum_{j=2}^{\infty} q^{-jk} \\
   &  = & \sum_{k=1}^{\infty} k P(k) \frac{1}{q^k(q^k-1)}.
\end{eqnarray*}
\end{lem}
\begin{proof}
For all $n \in \N$, we have that
\begin{eqnarray*}
n G(n) & = & \sum_{a \in G_n} \deg(a) \\
& = & \sum_{a \in G_n} \sum_{p^j | a} \deg(p) \\
& = & \sum_p \deg(p) \sum_{j=1}^{\infty} G(n-j\deg(p)) \\
& = & \sum_{k=1}^{\infty} P(k) k \sum_{j=1}^{\infty} G(n-jk).  
\end{eqnarray*}

If $n$ is sufficiently large, we may divide both sides by $G(n)$, 
and \A implies
\[ \sum_{k=1}^{\infty} P(k) k \sum_{j=2}^{\infty} \frac{G(n-jk)}{G(n)} \rightarrow C_3 \]
as $n \rightarrow \infty$ (note that the inner sum is actually finite), such that the lemma follows.
\end{proof}

\begin{sa}
\label{ControlError}
Suppose that $\lambda_n$ is bounded.
Let $r(n) = O(f(n))$ for a continuous nonincreasing function $f$ with $\lim_{x \rightarrow \infty} f(x) = 0$; 
let $F(x):= \int_{1}^{x} f(t) \, dt$. Suppose that $\int_1^{\infty} F(x)/x^2 \, dx$ converges. 
Then 
\begin{itemize}
\item (\ref{Pkkqkestimate}) holds with error term $O(F(n))$.
\item (\ref{Pkqkestimate2}) holds with error term $O(\int_n^{\infty} F(x)/x^2 \, dx)$.
\end{itemize}
\end{sa}
\begin{proof}
First note that $x f(x) = O(F(x))$.
In view of the previous theorem, we have to estimate
\begin{equation}
\label{diff}
\sum_{k=1}^n k P(k) \left(\frac{G(n-k)}{G(n)} - q^{-k}\right) = \sum_{k=1}^n \lambda_k \frac{r(n-k)-r(n)}{A+r(n)}.  
\end{equation}
such that our first assertion follows by reordering the sum $\sum_{k} r(n-k) = \sum_k r(k)$ and comparing it to an integral.

The second assertion follows by partial summation:
Let $S(t):= \sum_{k=1}^{\lfloor t \rfloor} \lambda_k $.
Let $s(t)$ defined by $S(t) = t + s(t) $; then $s(t) = O(F(t))$ by the first part. Let 
\[ I:=  \int_{1}^{\infty} \frac{s(t)}{t^2} \, dt.\]    
Then
\begin{eqnarray*}
\sum_{k=1}^n \lambda_k \cdot \frac{1}{k} & = & S(n) \cdot \frac{1}{n} + \int_1^n S(t) \frac{1}{t^2} \, dt \\   
& = & (n+s(n)) \cdot \frac{1}{n} + \int_1^n \frac{1}{t} \, dt + \int_{1}^n
\frac{s(t)}{t^2} \, dt \\
& = & 1  + O(F(n)/n) +  \ln(n) + I + O\left(\int_{n}^{\infty} \frac{F(t)}{t^2} \, dt\right)  
\end{eqnarray*}
Since $-F(n)/n = \int_n^{\infty} f(x)/x \, dx - \int_{n}^{\infty} F(x)/x^2 \, dx$, it follows that the error term is 
$O(\int_{n}^{\infty} F(x)/x^2 \, dx)$.   
(The main terms are, of course, those of (\ref{Pkqkestimate}), such that
$I + 1 = \gamma + \ln(A) - C_M $.) 
\end{proof}

It follows that Proposition \ref{TheoremErrorTerm} can be improved slightly.
\begin{cor}
The assertions of Proposition \ref{TheoremErrorTerm} hold under the weaker hypotheses
$P(n) = O(q^n/n)$, that $r(n) = O(1/n)$, and that $\sum_{n=1}^{\infty} |r(n)| $ converges.
\end{cor}
\begin{proof}
From our assumptions, it is clear that the right hand side of (\ref{diff}) is bounded in this case. 
The second assertion follows by partial summation as before (this time with $F(x)$ replaced by $1$), and the third one by exponentiation of the second and the definition of $C_M$.
\end{proof}

\section{A sum considered by Meissel}

The following, nearly forgotten result predates Mertens'
  1874 paper \cite{Mertens} containing the theorem named after him. See \cite{LindqvistPeetre2} for a proof and a full account of the historical background.
  
\begin{prop}[Meissel 1866]
For $\alpha \rightarrow 0+ $,
\[ \sum_{p \in \P} \frac{1}{p (\ln(p))^{\alpha}} = \frac{1}{\alpha} + C + O(\alpha), \]
where 
\[ C = \lim_{n \rightarrow \infty} \left(\sum_{p \leq n} \frac{1}{p} -
  \ln(\ln(n))\right). \]
\end{prop}

We show that this result, too, carries over to the case of additive arithmetical semigroups:
\begin{sa}
\label{Meisselsa}
Suppose $H(y) \rightarrow A$ for $ y \rightarrow 1/q$.
Let $C_1 :=  \ln(A) + \gamma - C_M $.
Let $\alpha > 0$. Let 
\[ S(x) := \sum_{k=1}^{\lfloor x \rfloor} \frac{P(k)}{q^k} \] 
and $s(x) := S(x) - \ln(x) - C_1$. Let 
\begin{equation}
\label{Jdef}
J(\alpha) := \int_1^{\infty} \frac{s(x)}{x^{\alpha+1}} \, dx
\end{equation}
Then 
\[ \sum_{k=1}^{\infty} \frac{P(k)}{q^k k^{\alpha}} = 1/\alpha + C_1 + \alpha J(\alpha)  \]
\end{sa}
\begin{proof}
By partial summation. 

\begin{eqnarray*}
\sum_{k=1}^n \frac{P(k)}{q^k k^{\alpha}} & = & S(n) \frac{1}{n^{\alpha}} +
\int_{1}^{n} S(x) \frac{\alpha}{x^{\alpha+1}} \, dx  \\  
& = & \frac{\ln(n) + C_1 + s(n)}{n^{\alpha}} + \alpha \int_{1}^{n}
\frac{\ln(x) + C_1 + s(x)}{x^{\alpha+1}} \, dx.
\end{eqnarray*}
Letting $n \rightarrow \infty$, the first summand on the right vanishes, and the summands of the integral are 
\[
\int_{1}^{\infty} \frac{\ln(x)}{x^{\alpha+1}} \, dx = \frac{1}{\alpha^2},
\]
\[ \int_1^{\infty} \frac{C_1}{x^{\alpha+1}} \, dx =  \frac{C_1}{\alpha}, \]
and (\ref{Jdef}), from which our result follows.
\end{proof}

To arrive at the same result as Meissel, we need $J(0)$ to exist. By Theorem~\ref{ControlError}, it suffices 
if $\int_t^{\infty} F(x)/x^2 dx = O(1/\ln(t)^{1+\epsilon})$ holds.
This finally gives us Meissel's theorem for additive arithmetical semigroups: 
\begin{cor}
If \A{} holds with error term $r(n) = O(1/\ln(n)^{2+\epsilon})$ and $P(n) = O(q^n/n)$, then
\[ \sum_{k=1}^{\infty} \frac{P(k)}{q^k k^{\alpha}} = 1/\alpha + C_1 + O(\alpha)  \]
as $\alpha \rightarrow 0$.
\end{cor}

\bibliography{arisemeng}

\small{
Authors's address: Stefan Wehmeier, Fakult\"at EIM, Universit\"at Paderborn, 33095 Paderborn, Germany. EMail: stefanw@math.upb.de.
}

\end{document}